### v1

# The hybrid matching of Hurwitz systems

Abstract.
In this paper we study planar hybrid systems composed by two stable linear systems, defined by Hurwitz matrices, in addition with a jump that can be a piecewise linear, a polynomial or an analytic function. We provide an explicit analytic necessary and sufficient condition for this class of hybrid systems to be asymptotically stable. We also prove the existence of limit cycles in this class of hybrid systems. Our results can be seen as generalizations of results already obtained in the literature. This was possible due to an embedding of piecewise smooth vector fields in a hybrid structure.
Key words and phrases:
Hybrid dynamical system; Global stability; Limit cycle.
2020 Mathematics Subject Classification:
34A38, 34D2 and 34A26.
1. Introduction
Consider a real autonomous system of differential equations
(1) $\dot{x}_{1}=P_{1}(x),\quad\dots,\quad\dot{x}_{n}=P_{n}(x),$
with $x=(x_{1},\dots,x_{n})\in\mathbb{R}^{n}$ and $P_{1},\dots,P_{n}\colon\mathbb{R}^{n}\to\mathbb{R}$ of class $C^{1}$, where the dot denotes the derivative with respect to the independent variable $t$ (time). As usual, we also identify system (1) with the vector field $X=(P_{1},\dots,P_{n})$.
We say that $x_{0}\in\mathbb{R}^{n}$ is a singularity of (1) if $X(x_{0})=0$. We say that a singularity $x_{0}$ is locally asymptotically stable if there is a neighborhood $U\subset\mathbb{R}^{n}$ of $x_{0}$ such that all orbits of (1) with initial points in $U$ tend to $x_{0}$ in forward time. The basin of attraction of $x_{0}$ is the largest $U$ satisfying this condition. If the basin of attraction is the entire $\mathbb{R}^{n}$, then we say that $x_{0}$ is globally asymptotically stable (GAS). After a translation if necessary, observe that we can assume that $x_{0}$ is the origin. If the origin is GAS, then by abuse of notation we say that $X$ is GAS.
The ability to determine the basin of attraction of a singularity is of great importance for applications of systems of ordinary differential equations (ODE). For example in evolutionary games (e.g. a game that models a conflict between different species of animals [31] or the level of corruption of a democratic society [1, 2]), an evolutionary stable strategy (ESS) represents a “uninvadable” state of the population, in the sense that small deviants behavior will eventually disappear under natural selection [15]. Under the language of ODE an ESS is represented by a locally asymptotically stable singularity [19, 27] and its basin of attraction determines how deviant a behavior can be and yet be tamed by natural selection. In particular if an ESS is GAS, then this mean that natural selection will eliminate any deviant behavior. For more information about ESS and its relation to ODE, we refer to [3, 4, 16] and the references therein.
Despite this importance, so far there is very few practical methods to determine if a given singularity is GAS. Among them one that stands out is the Markus-Yamabe condition, related with the following conjecture.
Conjecture 1 (Markus-Yamabe [26]).
Let $X=(P_{1},\dots,P_{n})$ be an autonomous $C^{1}$-vector field on $\mathbb{R}^{n}$ having a unique singularity at the origin. If the eigenvalues of the Jacobian matrix $DX(x)$ have negative real part for every $x\in\mathbb{R}^{n}$, then $X$ is GAS.
The conjecture was proved to be true for $n=2$ independently by Feßler [13], Glutsyuk [17] and Gutierrez [18]. On the other hand, Cima et al [10] proved that it is false for $n\geqslant 3$.
Although for $n=2$ the Markus-Yamabe condition is a sufficient condition for GAS, it is not necessary. In fact, consider the vector field $X=(P,Q)$ given by
$P(x,y)=-x+xy,\quad Q(x,y)=-y.$
Note that the origin is the unique singularity and that it is GAS, due to the existence of a global proper Lyapunov function $L\colon\mathbb{R}^{2}\to\mathbb{R}$ given by
$L(x,y)=\ln(1+x^{2})+y^{2}.$
Nevertheless, the eigenvalues of $DX(x,y)$ are given by $\lambda_{1}=-1$ and $\lambda_{2}=-1+y$ and thus $X$ does not satisfy the Markus-Yamabe condition.
To this end we shall say that a planar $C^{1}$-vector field satisfying the Markus-Yamabe condition is a Markus-Yamabe vector field (MY-vector field). In particular observe that if $X$ is a linear MY-vector field, then it is given by
$X(x,y)=A\left(\begin{array}[]{c}x\\ y\end{array}\right),$
where $A$ is a $2\times 2$ matrix with eigenvalues having negative real part. Since such matrices are known as Hurwitz matrices [12], we say that such a vector field is a Hurwitz vector field.
This lack of practical conditions for GAS is also a problem for other types of vector fields, such as the planar piecewise smooth vector fields. Briefly, we recall that a planar piecewise smooth vector field is a tuple $\mathcal{X}=(X^{+},X^{-};\Sigma)$ such that $\Sigma=h^{-1}(\{0\})$, where $h\colon\mathbb{R}^{2}\to\mathbb{R}$ is a continuous function and $X^{\pm}$ are planar $C^{1}$-vector fields defined in a neighborhood of
$\Sigma^{\pm}=\{(x,y)\in\mathbb{R}^{2}\colon\pm h(x,y)\geqslant 0\}.$
Let $q\in\mathbb{R}^{2}$. If $q\in\Sigma^{\pm}\setminus\Sigma$, then the local trajectory of $\mathcal{X}$ at $q$ is given by the local trajectory of $X^{\pm}$ at $q$. If $q\in\Sigma$, then the local trajectory can be classified as crossing, sliding or escaping. Moreover, $q$ can also be a new type of singularity, known as tangential singularity. In this paper we only deal with crossing points. For the definition of the other types of trajectories and for more details on piecewise smooth vector fields, we refer to [11] and the references therein.
We say that $q\in\Sigma$ is a crossing point of $\mathcal{X}$ if
(2) $X^{+}(q),\nabla h(q)\right>\cdot\left<X^{-}(q),\nabla h(q)\right>>0,$
[…]
As far as we know the first approach on the characterization of global asymptotic stability for piecewise smooth vector fields is due Freire to et al [14], where they proved that the continuous matching of two Hurwitz vector fields is GAS. More precisely, if $\mathcal{X}=(X^{+},X^{-};\Sigma)$ is a planar piecewise linear vector field satisfying the following hypotheses:
  1. $(H_{1})$
    $X^{\pm}$ are Hurwitz vector fields;
  2. $(H_{2})$
    $\Sigma$ is a straight line containing the origin;
  3. $(H_{3})$
    $X^{+}|_{\Sigma}=X^{-}|_{\Sigma}$;
then it is GAS. Braga et al [9] extended this result by proving that if we replace $H_{3}$ by:
  1. $(H_{3}^{\prime})$
    The points on $\Sigma\setminus\{0\}$ are of crossing type;
then $\mathcal{X}$ is also GAS. However, later it was proved that further generalizations does not maintain the global stability. More precisely, consider the following generalizations for $H_{1}$ and $H_{2}$:
  1. $(H_{1}^{\prime})$
    $X^{\pm}$ are GAS;
  2. $(H_{1}^{\prime\prime})$
    $X^{\pm}$ are MY-vector fields;
  3. $(H_{2}^{\prime})$
    $\Sigma$ is a polygonal line containing the origin.
It follows from [9, 7, 8] that none of the following conditions
$H_{1}^{\prime}\cap H_{2}\cap H_{3}^{\prime},\quad H_{1}^{\prime\prime}\cap H_{2}\cap H_{3}^{\prime},\quad H_{1}\cap H_{2}^{\prime}\cap H_{3}^{\prime},$
imply the global asymptotically stability for $\mathcal{X}$.
In this paper we generalize these results in a different way. We embed the piecewise smooth vector field $\mathcal{X}$ in a hybrid structure, defining a hybrid system. Briefly, a hybrid dynamical system is a system whose dynamics are governed both by continuous and discrete laws. That is, a system that can both flow and jump [23].
Within this structure we are able to unify the hypotheses
$H_{1}\cap H_{2}\cap H_{3}^{\prime}\quad\text{and}\quad H_{1}\cap H_{2}^{\prime}\cap H_{3}^{\prime},$
understanding precisely when this new hybrid system is GAS. Moreover new bifurcations are now possible, such as limit cycles.
The paper is organized as follows. In Section 2 we provide a brief introduction to hybrid systems. The statement of the main result is postponed to Section 3. In Section 4 we have some preliminary and technical results used in the proof of the main result, whose proof is given in Section 5. Finally in Section 6 we have a conclusion and some further thoughts.
2. Hybrid systems
Before we state what we mean by a hybrid system in this paper, we present one of the usual examples of the field, the bouncing ball model.
Example 1 (Example $1$ of [23]).
Consider the vertical motion of a ball dropped (or tossed) from an initial height $h>0$, with initial velocity $v\in\mathbb{R}$ (here negative velocity means downwards, following gravity, while positive velocity means that the ball was tossed upwards). If the ball is under the acceleration of constant gravity $g>0$, then for $h>0$ the state of the ball is governed by the system of differential equations
(3) $\dot{h}=v,\quad\dot{v}=-g.$
Therefore given any initial condition $q=(h_{0},v_{0})$, $h_{0}>0$, it follows from (3) that the ball reaches the ground $h=0$ after a finite amount of time $t_{0}>0$, with velocity $v(t_{0})<0$. See Figure 1.
[…]
Figure 1. Illustration of an orbit of the bouncing ball model. For simplicity, we interchanged the coordinate axes.
At this instant, the ball bounces back upwards and its velocity undergoes an instantaneous change $v(t_{0})\mapsto-\rho v(t_{0})$, modeled by an inelastic collision with dissipate factor $10<\rho<$ 1. That is, at $h=0$ our system undergoes a jump (or reset) $\varphi$ given by $\varphi(0,v)=(0,-\rho v)$. Roughly speaking, the reset map represents the instantaneous loss of energy that the system suffers when hitting the ground. Observe that any orbit converges to the origin, which represents a ball standing still in the ground.
For more details on the bouncing ball model, we refer to [30, Section $2.2.3$] and [6, Section $1.2$]. For a different example modeling a type of pinball machine with curved surface, we refer to [29, Example $2$]. For a recent survey in the field, we refer to [5].
The notion of a hybrid dynamical system is relatively new. Following Schaft and Schumacher [30, Section $1.1$], we quote:
[$\dots$] the area of hybrid systems is still in its infancy and a general theory of hybrid systems seems premature. More inherently, hybrid systems is such a wide notion that sticking to a single definition shall be too restrictive. [$\dots$] Another difficulty in discussing hybrid systems is that various scientific communities with their own approaches have contributed (and are still contributing) to the area. At least the following three communities can be distinguished. [The] computer science community [$\dots$] [the] modeling and simulation community [$\dots$] [and the] systems and control community.
Nevertheless, the field is expanding and in particular it has been receiving more attention in recent years from the community of Qualitative Theory of Differential Equations and Dynamical Systems. See for example the recent works on Melnikov method [21, 22], limit cycles [25], topological horseshoes [32], chaos [23] and polycycles [29].
One fact that draw attention to the hybrid systems is that a piecewise smooth vector field can be embedded in a hybrid structure, allowing the existence of limit cycles that otherwise would not be possible. For example Llibre and Teixeira [24] proved that a planar piecewise linear vector field given by two centers cannot have limit cycles. However it follows from [20, 23] that if we embed such systems in a hybrid structure, then we have the existence of at least one limit cycle.
As anticipated in the Introduction, in this paper we will embedded the the piecewise smooth vector field $\mathcal{X}=(X^{+},X^{-};\Sigma$ in a hybrid structure. This embedding will allow us to draw a precise sufficient and necessary condition for global stability and moreover will point the way for the bifurcation of a limit cycle that would not be possible without the hybrid framework.
We now establish what we mean by a planar hybrid system in this paper. Similarly to the definition of a piecewise vector field, a hybrid system is a tuple $\mathfrak{X}=(X^{+},X^{-};\Sigma;\varphi)$ such that $\Sigma=h^{-1}(\{0\})$, where $h\colon\mathbb{R}^{2}\to\mathbb{R}$ is a continuous function; $X^{\pm}$ are planar $C^{1}$-vector fields defined in a neighborhood of
$\Sigma^{\pm}=\{(x,y)\in\mathbb{R}^{2}\colon\pm h(x,y)\geqslant 0\},$
and a map $\varphi\colon\Sigma\to\Sigma$ known as jump (or reset map).
In this paper $X^{\pm}$ are Hurwitz vector fields, $\Sigma=\Sigma_{\rho}$ is the zero locus of the function $h_{\rho}\colon\mathbb{R}^{2}\to\mathbb{R}$ given by
$h_{\rho}(x,y)=\left\{\begin{array}[]{ll}y,&\text{if }x\leqslant 0,\\ y-\rho x,&\text{if }x\geqslant 0,\end{array}\right.$
where $\rho\in\mathbb{R}_{\geqslant 0}$. That is $\Sigma_{\rho}$ is the polygonal line (also known as broken line) given by $\Sigma_{\rho}=\Sigma^{1}\cup\Sigma^{2}_{\rho}$, where
(4) $\Sigma^{1}=\{(x,y)\in\mathbb{R}^{2}\colon x\leqslant 0,y=0\},\quad\Sigma^{2}_{\rho}=\{(x,y)\in\mathbb{R}^{2}\colon x\geqslant 0,y=\rho x\}.$
We shall also suppose that $\mathfrak{X}$ has the crossing property, i.e. we suppose that
(5) $X^{+}(q),\nabla h_{\rho}(q)\right>\cdot\left<X^{-}(q),\nabla h_{\rho}(q)\right>>0,$
for every $q\in\Sigma_{\rho}\setminus\{(0,0)\}$. In particular we recall that the geometric interpretation of (5) is that $X^{\pm}(q)$ are transversal to $\Sigma$ at $q$ and that both vectors $X^{+}(q)$ and $X^{-}(q)$ points to the same direction relatively to $\nabla h_{\rho}(q)$. For hybrid systems without the crossing property and thus allowing other dynamics such as sliding (or sticking), we refer to [30, Section $6.2.3$] and [6, p. $82$].
The jump $\varphi=\varphi_{\rho}\colon\Sigma_{\rho}\to\Sigma_{\rho}$ is given by
(6) $\varphi_{\rho}(x,y)=\left\{\begin{array}[]{ll}(-a|x|^{r},0),&\text{if }x\leqslant 0,\vspace{0.2cm}\\ (bx^{s},\rho bx^{s}),&\text{if }x\geqslant 0,\end{array}\right.$
with $a$, $b$, $r$, $s\in\mathbb{R}_{>0}$. Observe that $\varphi_{\rho}$ has inverse given by
(7) $\varphi_{\rho}^{-1}(x,y)=\left\{\begin{array}[]{ll}\bigl{(}-\bigl{|}a^{-1}x\bigr{|}^{\frac{1}{r}},0\bigr{)},&\text{if }x\leqslant 0,\vspace{0.2cm}\\ \bigl{(}b^{-\frac{1}{s}}x^{\frac{1}{s}},\rho b^{-\frac{1}{s}}x^{\frac{1}{s}}\bigr{)},&\text{if }x\geqslant 0,\end{array}\right.$
for each $\rho\in\mathbb{R}_{\rho\geqslant 0}$. Moreover $\Sigma^{1}$ and $\Sigma^{2}_{\rho}$ are invariant by $\varphi_{\rho}$, i.e. $\varphi_{\rho}(\Sigma^{1})=\Sigma^{1}$ and $\varphi_{\rho}(\Sigma^{2}_{\rho})=\Sigma^{2}_{\rho}$. See Figure 2.
[…]
Figure 2. Illustration of $\varphi_{\rho}$ and $\mathfrak{X}$. The origin $\mathcal{O}$ is interpreted as a singularity. By abuse of notation we drew two copies of $\Sigma_{\rho}$.
The dynamics of $\mathfrak{X}=(X^{+},X^{-};\Sigma;\varphi)$ works as follows. Given $q_{0}\in\mathbb{R}^{2}\setminus\Sigma$, the local trajectory of $\mathfrak{X}$ at $q_{0}$ is given by the local trajectory of $X^{\pm}$. If this trajectory never intersects $\Sigma_{\rho}$, then we are done. If it does intersect $\Sigma$ at a point $q_{1}$, then we apply the jump obtaining a point $q_{2}=\varphi_{\rho}(q_{1})$. It follows from (5) that the local trajectories of $X^{\pm}$ at $q_{2}$ are transversal to $\Sigma$ and agree on orientation. Hence we are able to follow exactly one of such trajectories leaving $\Sigma$. The process now repeats. See Figure 2.
In simple words, the dynamic of the hybrid system $\mathfrak{X}=(X^{+},X^{-};\Sigma_{\rho};\varphi)$ is similar to the dynamics of the piecewise vector field $\mathcal{X}=(X^{+},X^{-};\Sigma_{\rho})$, with the exception that when hitting the switching set $\Sigma$ we apply the jump before crossing it. In particular observe from (5) that if $a=b=r=s=1$, then $\varphi_{\rho}$ reduces to the identity map and thus the hybrid system becomes a piecewise vector field.
Similarly to the fact that switching between $X^{+}$ and $X^{-}$ has the physic interpretation of a particle switching from one medium to another (e.g. the refraction of light when passing from air to water), the jump $\varphi_{\rho}$ represents the instantaneous gain or loss of energy that such a particle may suffer when transitioning between the mediums.
3. Statement of the main result
We now provide the definitions for a precise statement of our main result. Let $\mathfrak{X}=(X^{+},X^{-};\Sigma_{\rho};\varphi_{\rho})$ be a hybrid system satisfying the following hypotheses:
  1. $(H_{1})$
    $X^{\pm}$ are Hurwitz vector fields;
  2. $(H_{2}^{\prime})$
    $\Sigma_{\rho}=\Sigma^{1}\cup\Sigma^{2}_{\rho}$ is the polygonal straight line given by (4);
  3. $(H_{3}^{\prime})$
    $\mathfrak{X}$ has the crossing property (5).
It follows from $H_{1}$ that $X^{\pm}$ can only have the following types of singularities at the origin:
  1. (i)
    An attracting node with distinct eigenvalues, denoted by $N_{1}$;
  2. (ii)
    An attracting non-diagonalizable node, denoted by $N_{2}$;
  3. (iii)
    An attracting focus, denoted by $F$.
In particular, note that if $X^{\pm}$ have an attracting star node (i.e. a diagonalizable node with equal eigenvalues), then it does satisfy $H_{1}$. In our main result we provide a complete characterization of the dynamics of $\mathfrak{X}$.
Theorem A.
Let $\mathfrak{X}=(X^{+},X^{-};\Sigma_{\rho};\varphi_{\rho})$ be a hybrid system satisfying hypotheses $H_{1}$, $H_{2}^{\prime}$ and $H_{3}^{\prime}$. Then the following statements hold.
  1. (a)
    If $X^{+}$ or $X^{-}$ is not of type $F$, then $\mathfrak{X}$ is GAS.
  2. (b)
    If $X^{+}$ and $X^{-}$ are both of type $F$, then exactly one of the following statements hold.
    1. (i)
      The origin is GAS.
    2. (ii)
      The origin is globally asymptotically unstable.
    3. (iii)
      The origin is a global center.
    4. (iv)
      $\mathfrak{X}$ has a unique limit cycle which is hyperbolic.
Moreover, all statements occur and there is an explicit analytic characterization of it.
4. Preliminary results
In this section we provide some technical results on global normal forms for $\mathfrak{X}$ and on the dynamics of Hurwitz linear vector fields of type $N_{1}$, $N_{2}$ and $F$.
4.1. Global normal form
Lemma 1.
Let $\mathfrak{X}=(X^{+},X^{-};\Sigma_{\rho};\varphi_{\rho})$ be a hybrid system satisfying hypotheses $H_{1}$, $H_{2}^{\prime}$ and $H_{3}^{\prime}$. Let also $B^{\pm}$ be the Hurwitz matrices such that
$X^{\pm}(x,y)=B^{\pm}\left(\begin{array}[]{c}x\\ y\end{array}\right).$
Then there is a topological equivalence between the piecewise linear systems $(X^{+},X^{-};\Sigma_{\rho})$ and $(Y^{+},Y^{-};\Sigma_{\rho})$, where $Y^{\pm}$ are given by
$Y^{\pm}(x,y)=A^{\pm}\left(\begin{array}[]{c}x\\ y\end{array}\right),\quad A=\left(\begin{array}[]{cc}\sigma^{\pm}&\delta^{\pm}\\ 1&0\end{array}\right),$
with $\sigma^{\pm}=\operatorname{tr}B^{\pm}<0$ and $\delta^{\pm}=-\det B^{\pm}<0$.
Proof.
If we let
$B^{\pm}=\left(\begin{array}[]{cc}b_{11}^{\pm}&b_{12}^{\pm}\vspace{0.2cm}\\ b_{21}^{\pm}&b_{22}^{\pm}\end{array}\right),$
then the crossing property (5) restricted at $\Sigma^{2}_{\rho}$ writes
$\bigl{(}b_{21}^{+}+(b_{22}^{+}-b_{11}^{+})\rho-b_{12}^{+}\rho^{2}\bigr{)}\bigl{(}b_{21}^{-}+(b_{22}^{-}-b_{11}^{-})\rho-b_{12}^{-}\rho^{2}\bigr{)}x^{2}>0.$
In particular, we obtain
(8) $\eta^{\pm}:=b_{21}^{\pm}+(b_{22}^{\pm}-b_{11}^{\pm})\rho-b_{12}^{\pm}\rho^{2}\neq 0.$
Consider the matrices
$C^{\pm}=\left(\begin{array}[]{cc}c_{11}^{\pm}&c_{12}^{\pm}\vspace{0.2cm}\\ c_{21}^{\pm}&c_{22}^{\pm}\end{array}\right),$
where
$\begin{array}[]{cc}\displaystyle c_{11}^{\pm}=1-\frac{(\delta^{\pm}-b_{12}^{\pm})\rho+b_{22}}{\eta^{\pm}}\rho,&\displaystyle c_{12}^{\pm}=\frac{(\delta^{\pm}-b_{12}^{\pm})\rho+b_{22}}{\eta^{\pm}},\vspace{0.2cm}\\ \displaystyle c_{21}^{\pm}=\rho+\frac{b_{12}^{\pm}\rho^{2}+b_{11}^{\pm}\rho-1}{\eta^{\pm}}\rho,&\displaystyle c_{22}^{\pm}=-\frac{b_{12}^{\pm}\rho^{2}+b_{11}^{\pm}\rho-1}{\eta^{\pm}}.\end{array}$
It follows from (8) that $C^{\pm}$ are well defined. Moreover we observe that
$\det C^{\pm}=-\frac{\delta^{\pm}\rho^{2}+\sigma^{\pm}\rho-1}{\eta^{\pm}}.$
Hence it follows from (8), $\delta^{\pm}<0$, $\sigma^{\pm}<0$ and $\rho\geqslant 0$ that $\det C^{\pm}\neq 0$ and thus $C^{\pm}$ is a linear change of variables. From straightforward calculations one can see that
$A^{\pm}=C^{\pm}B^{\pm}(C^{\pm})^{-1}\quad\text{and}\quad C^{\pm}\left(\begin{array}[]{c}x\\ y\end{array}\right)=\left(\begin{array}[]{c}x\\ y\end{array}\right),\quad\text{if}\quad(x,y)\in\Sigma_{\rho}.$
Therefore if we define the map $H\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$ by
$H(x,y)=\left\{\begin{array}[]{ll}C^{+}\left(\begin{array}[]{c}x\\ y\end{array}\right),&\text{if }(x,y)\in\Sigma^{+},\vspace{0.2cm}\\ C^{-}\left(\begin{array}[]{c}x\\ y\end{array}\right),&\text{if }(x,y)\in\Sigma^{-},\end{array}\right.$
then it is a well defined topological equivalence between the piecewise systems $(X^{+},X^{-};\Sigma_{\rho})$ and $(Y^{+},Y^{-};\Sigma_{\rho})$. ∎
For simplicity throughout the remaining of this section we state our results only for $A^{+}$. Similar results can be proved for $A^{-}$.
4.2. Hurwitz vector fields of type $N_{1}$
Suppose that the origin is an attracting node of $X^{+}$ with distinct eigenvalues given by
$r_{1}^{+}=\frac{\sigma^{+}}{2}+\frac{\sqrt{(\sigma^{+})^{2}+4\delta^{+}}}{2},\quad r_{2}^{+}=\frac{\sigma^{+}}{2}-\frac{\sqrt{(\sigma^{+})^{2}+4\delta^{+}}}{2}.$
In particular, observe that $r_{2}^{+}<r_{1}^{+}<0$. Following [9] we define the sets
$\begin{array}[]{l}R_{1}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x>r_{1}^{+}y\},\vspace{0.2cm}\\ E_{1}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x=r_{1}^{+}y\},\vspace{0.2cm}\\ R_{2}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon r_{2}^{+}y<x<r_{1}^{+}y\},\vspace{0.2cm}\\ E_{2}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x=r_{2}^{+}y\},\vspace{0.2cm}\\ R_{3}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x<r_{2}^{+}y\},\end{array}$
see Figure 3.
[…]
Figure 3. Illustration of two Hurwitz vector fields $X^{\pm}$ of type $N_{1}$. The curves in red (resp. blue) are the solutions of $X^{+}$ (resp. $X^{-}$). For interpretation of the references to color in this figure legend, the reader is referred to the web version of this article.
Given $q\in\Sigma_{\rho}^{+}$, let $X^{+}(t,q)$ denote the solution of $X^{+}$ with initial condition $X^{+}(0,q)=0$. The following lemma follows similarly to [9, Lemma $5$].
Lemma 2.
Let $X^{+}$ be a Hurwitz vector field of type $N_{1}$ defined in $\Sigma_{\rho}^{+}$. Then the following statements hold (see Figure 3).
  1. (a)
    If $q\in E_{1}^{+}\cup E_{2}^{+}$, then $X^{+}(t,q)\to\mathcal{O}$ as $t\to+\infty$.
  2. (b)
    If $q\in R_{1}^{+}$, then
    1. (i)
      $X^{+}(t,q)\to\mathcal{O}$ as $t\to+\infty$;
    2. (ii)
      there is $t^{1}_{q}<0$ such that $X^{+}(t^{1}_{q},q)\in\Sigma_{\rho}$.
  3. (c)
    If $q\in R_{2}^{+}$, then $X^{+}(t,q)\to\mathcal{O}$ as $t\to+\infty$.
  4. (d)
    If $q\in R_{3}^{+}$, then there is $t^{2}_{q}>0$ such that $X^{+}(t^{2}_{q},q)\in\Sigma_{\rho}$.
4.3. Hurwitz vector fields of type $N_{2}$
Suppose that the origin is an attracting non-diagonalizable node of $X^{+}$ with repeated eigenvalue given by
$r^{+}=\frac{\sigma^{+}}{2}<0.$
Following [9] we define the sets
$\begin{array}[]{l}S_{1}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x>r^{+}y\},\vspace{0.2cm}\\ E^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x=r^{+}y\},\vspace{0.2cm}\\ S_{2}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x<r^{+}y\},\end{array}$
see Figure 4.
[…]
Figure 4. Illustration of two Hurwitz vector fields $X^{\pm}$ of type $N_{2}$. The curves in red (resp. blue) are the solutions of $X^{+}$ (resp. $X^{-}$). For interpretation of the references to color in this figure legend, the reader is referred to the web version of this article.
The following lemma follows from [9, Lemma $6$].
Lemma 3.
Let $X^{+}$ be a Hurwitz vector field of type $N_{2}$ defined in $\Sigma_{\rho}^{+}$. Then the following statements hold (see Figure 4).
  1. (a)
    If $q\in E^{+}$, then $X^{+}(t,q)\to\mathcal{O}$ as $t\to+\infty$.
  2. (b)
    If $q\in S_{1}^{+}$, then
    1. (i)
      $X^{+}(t,q)\to\mathcal{O}$ as $t\to+\infty$;
    2. (ii)
      there is $t^{1}_{q}<0$ such that $X^{+}(t^{1}_{q},q)\in\Sigma_{\rho}$.
  3. (c)
    If $q\in S_{2}^{+}$, then there is $t^{2}_{q}>0$ such that $X^{+}(t^{2}_{q},q)\in\Sigma_{\rho}$.
4.4. Hurwitz vector fields of type $F$
Suppose that the origin is an attracting focus of $X^{+}$ with eigenvalues given by
$r_{1}^{+}=\lambda^{+}+i\mu_{+},\quad r_{2}^{+}=\lambda^{+}-i\mu_{+},$
where
$\lambda^{+}=\frac{\sigma^{+}}{2}<0,\quad\mu^{+}=\frac{\sqrt{|(\sigma^{+})^{2}+4\delta^{+}|}}{2}>0,$
and $i$ is the imaginary unit satisfying $i^{2}=-1$. The following lemma follows from [9, Lemma $7$].
Lemma 4.
Let $X^{+}$ be a Hurwitz vector field of type $F$ defined in $\Sigma_{\rho}^{+}$. Then for every $q\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}$ there are $t^{1}_{q}<0<t^{2}_{q}$ such that $X^{+}(t^{1}_{q},q)\in\Sigma_{\rho}$ and $X^{+}(t^{2}_{q},q)\in\Sigma_{\rho}$ (see Figure 5).
[…]
Figure 5. Illustration of two Hurwitz vector fields $X^{\pm}$ of type $F$. The curves in red (resp. blue) are the solutions of $X^{+}$ (resp. $X^{-}$). For interpretation of the references to color in this figure legend, the reader is referred to the web version of this article.
To end this section, we observe that since $X^{+}$ is linear we can compute explicitly its solution, see [28, Chapter $1$]. Therefore it is not hard to see that if $X^{+}$ is a Hurwitz vector field of type $F$ with the normal form given by Lemma 1, then its solution
$X^{+}(t;x,y)=\bigl{(}x^{+}(t;x,y),y^{+}(t;x,y)\bigr{)}$
is given by
(9) $\begin{array}[]{l}\displaystyle x^{+}(t;x,y)=e^{\lambda^{+}t}\left(x\cos(\mu^{+}t)+\frac{1}{\mu^{+}}(\lambda^{+}x-r_{1}^{+}r_{2}^{+}y)\sin(\mu^{+}t)\right),\vspace{0.2cm}\\ \displaystyle y^{+}(t;x,y)=e^{\lambda^{+}t}\left(y\cos(\mu^{+}t)+\frac{1}{\mu^{+}}(x-\lambda^{+}y)\sin(\mu^{+}t)\right).\end{array}$
5. Proof of Theorem A
Proof of Theorem A.
It follows from Lemmas 2, 3 and 4 that if $X^{+}$ or $X^{-}$ is not of type $F$, then $\mathfrak{X}$ is GAS. More precisely, if we suppose for example that $X^{+}$ is of type $N_{1}$ and $X^{-}$ is of type $N_{2}$ (i.e. $N_{1}-N_{2}$), then it is clear that every orbit of $\mathfrak{X}$ will eventually go to the origin. See Figures 3 (in red) and 4 (in blue). The cases $N_{1}-N_{1}$, $N_{1}-F$, $N_{2}-N_{2}$, $N_{2}-F$ and their symmetric (e.g. $F-N_{1}$) follow similarly. This proves statement $(a)$.
We now focus on statement $(b)$. In particular, from now on we suppose that both $X^{\pm}$ are of type $F$. Consider the displacement map $\Delta\colon\mathbb{R}_{>0}\to\mathbb{R}$ given by
(10) $\Delta(x)=x^{-}\bigl{(}-t^{-};\varphi^{-1}_{\rho}(x,\rho x)\bigr{)}-\varphi_{\rho}\bigl{(}x^{+}(t^{+};x,\rho x)\bigr{)},$
where $t^{\pm}>0$ are such that
(11) $y^{+}(t^{+};x,\rho x)=0,\quad y^{-}\bigl{(}-t^{-};\varphi^{-1}_{\rho}(x,\rho x)\bigr{)}=0.$
See Figure 6.
[…]
Figure 6. Illustration of the displacement map $\Delta$.
From (6) and (9) we have
(12) $\varphi_{\rho}\bigl{(}x^{+}(t^{+};x,\rho x)\bigr{)}=-ae^{\lambda^{+}t^{+}r}|\cos(\mu^{+}t^{+})+\Phi^{+}\sin(\mu^{+}t^{+})|^{r}x^{r},$
where $\Phi^{\pm}=(\lambda^{\pm}-r_{1}^{\pm}r_{2}^{\pm}\rho)/\mu^{\pm}$. On the other hand, from (7) and (9) we obtain
(13) $x^{-}\bigl{(}-t^{-};\varphi^{-1}_{\rho}(x,\rho x)\bigr{)}=e^{-\lambda^{-}t^{-}}\bigl{(}\cos(\mu^{-}t^{-})-\Phi^{-}\sin(\mu^{-}t^{-})\bigr{)}b^{-\frac{1}{s}}x^{\frac{1}{s}}.$
Therefore it follows from (10), (12) and (13) that $\Delta(x)<0$ if and only if
(14) $ab^{\frac{1}{s}}x^{r}<e^{|\lambda^{-}t^{-}+\lambda^{+}t^{+}r|}\frac{|\cos(\mu^{-}t^{-})-\Phi^{-}\sin(\mu^{-}t^{-})|}{|\cos(\mu^{+}t^{+})+\Phi^{+}\sin(\mu^{+}t^{+})|^{r}}x^{\frac{1}{s}}.$
From (14) we can deduce statements $(i)$, $(ii)$, $(iii)$ and $(iv)$. More precisely, if $r=1/s$ then we can factor $x$ out of (14) and thus we obtain that the origin is globally asymptotically stable (resp. globally asymptotically unstable) if (14) is satisfied (resp. is the opposite inequality is satisfied). If the equality holds, then the origin is a global center.
If $r\neq 1/s$ then we can divide (14) either by $x^{r}$ or $x^{\frac{1}{s}}$ (whichever is smaller) and thus obtain that there is exactly one $x_{0}>0$ satisfying the equality in (14). This $x_{0}$ represents a limit cycle. Differentiating $\Delta(x)$ in this case one can see that $\Delta^{\prime}(x_{0})\neq 0$ and thus the limit cycle is hyperbolic. This finishes the proof of statement $(b)$. ∎
Remark 1 (The case $\rho=0$).
Solving (11) we obtain,
(15) $t^{\pm}(\rho)=\frac{1}{\mu^{\pm}}\left(\pm\arctan\left(\frac{\mu^{\pm}}{\lambda^{\pm}\rho-1}\rho\right)+\pi\right).$
In particular we have $t^{\pm}(0)=\pi/\mu^{\pm}$ and thus if we replace $\rho=0$ at (14) we obtain,
$ab^{\frac{1}{s}}x^{r}<e^{\bigl{|}\frac{\lambda^{-}}{\mu^{-}}+\frac{\lambda^{+}}{\mu^{+}}r\bigr{|}\pi}x^{\frac{1}{s}}.$
Observe that if $a=b=r=s=1$, then the above inequality is satisfied regardless of $\lambda^{\pm}$ and $\mu^{\pm}$, i.e. regardless of $X^{\pm}$. This in addition with Theorem A$(a)$ ensures that if the jump $\varphi_{\rho}$ is reduced to the identity map (i.e. if the hybrid system $\mathfrak{X}$ is reduced to a piecewise linear vector field) and $\Sigma_{\rho}$ is reduced to a straight line, then $\mathfrak{X}$ is GAS. This is precisely the main result of [9].
Remark 2 (The limit case $\rho\to+\infty$).
From (15) we have that,
$\lim\limits_{\rho\to+\infty}t^{\pm}(\rho)=\frac{1}{\mu^{\pm}}\left(\pm\arctan\left(\frac{\mu^{\pm}}{\lambda^{\pm}}\right)+\pi\right).$
This in addition with the following well-known properties of the trigonometric functions
$\begin{array}[]{ll}\displaystyle\cos(x+\pi)=-\cos x,&\displaystyle\sin(x+\pi)=-\sin(x),\vspace{0.2cm}\\ \displaystyle\cos(\arctan x)=\frac{1}{\sqrt{1+x^{2}}},&\displaystyle\sin(\arctan x)=\frac{x}{\sqrt{1+x^{2}}},\end{array}$
and some straightforward calculations, allow us to conclude that
(16) $\lim\limits_{\rho\to+\infty}\frac{|\cos(\mu^{-}t^{-})-\Phi^{-}\sin(\mu^{-}t^{-})|}{|\cos(\mu^{+}t^{+})+\Phi^{+}\sin(\mu^{+}t^{+})|^{r}}=\left\{\begin{array}[]{ll}0,&\text{if }r>1,\vspace{0.2cm}\\ +\infty,&\text{if }r<1,\vspace{0.2cm}\\ \displaystyle\sqrt{\frac{(\lambda^{-})^{2}+(\mu^{-})^{2}}{(\lambda^{+})^{2}+(\mu^{+})^{2}}},&\text{if }r=1.\end{array}\right.$
Since the above limit is the only factor of (14) that depends on $\rho$, we conclude that if $r\neq 1$ then the right-hand side of (14) can either explode or collapse to zero as $\rho\to+\infty$. In particular, if we assume that $1/s=r$ (and thus factoring $x$ out of (14)), then it follows from (16) that the stability of the origin can reverse as $\rho\to+\infty$.
We finish this section with an example that shows that even if $\Sigma_{\rho}$ a straight line (i.e. $\rho=0$) and $X^{\pm}$ has a continuous match at $\Sigma$, the hybrid system $\mathfrak{X}$ can have a limit cycle. We observe that the existence of such a limit cycle was proved impossible in the previous piecewise framework [9, 14].
Example 2.
Consider the hybrid system $\mathfrak{X}=(X^{+},X^{-};\Sigma;\varphi)$ obtained by replacing
(17) $a=b=1,\quad r=s=3,\quad\rho=0,\quad\lambda^{\pm}=-1,\quad\mu^{\pm}=1,$
in the proof of Theorem A. Using the global normal form provided by Lemma 1, we have that $\mathfrak{X}$ is given by
$\begin{array}[]{l}X^{\pm}(x,y)=(-2x-2y,x),\vspace{0.2cm}\\ \Sigma=\{(x,y)\in\mathbb{R}^{2}\colon y=0\},\vspace{0.2cm}\\ \varphi(x,0)=(x^{3},0).\end{array}$
Replacing (17) into (12), (13) and (15) we obtain
(18) $\Delta(x)=-e^{\pi}x^{\frac{1}{3}}+e^{-3\pi}x^{3},$
where we recall that $\Delta\colon\mathbb{R}_{>0}\to\mathbb{R}$ is the displacement map given by (10). Solving (18) we obtain that $x_{0}=e^{\frac{3}{2}\pi}$ is the unique solution of $\Delta(x)=0$. Hence, $\mathfrak{X}$ has a unique limit cycle $\gamma$. Differentiating (18) we obtain
$\Delta^{\prime}(x)=-\frac{1}{3}e^{\pi}x^{-\frac{2}{3}}+3e^{-3\pi}x^{2}.$
Since $\Delta^{\prime}(x_{0})=8/3$, we have that $\gamma$ is hyperbolic and unstable. In particular, from the uniqueness of $\gamma$ we have that the origin is locally asymptotically stable. See Figure 7.
[…]
Figure 7. Illustration of $\mathfrak{X}$. Here blue denotes stable and red denotes unstable. For interpretation of the references to color in this figure legend, the reader is referred to the web version of this article.
6. Conclusion and further thoughts
The main inspiration for this work is the sequence of papers [9, 7, 8] in which is studied the crossing matching of different types of globally asymptotically stable vector fields. Such matching sometimes result in a globally asymptotically stable piecewise vector field, but not always.
By embedding these piecewise vector fields in a hybrid structure we were able to connect such results in a unique formulae that encapsulates the global dynamics of this new hybrid system. In particular, such formulae shows precisely when the origin is GAS and also indicates the bifurcation of a limit cycle that was previously impossible in the piecewise framework.
Acknowledgments
The first author is partially supported by Fundação de Amparo à Pesquisa do Estado de Minas Gerais, Brazil, grant APQ-02153-23; by São Paulo Research Foundation (FAPESP), Brazil, grant 2019/07316–0; and by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Brazil, grant 311921/2020-5.
The second author is partially supported by São Paulo Research Foundation (FAPESP), Brazil, grant 2021/01799-9.
The second author is grateful for the hospitality of Universidade Federal de Itajubá (UNIFEI), where this study was developed.

### v2

@@
 If $q\in\Sigma$, then the local trajectory can be classified as crossing, sliding or escaping.
-Moreover, $q$ can also be a new type of singularity, known as tangential singularity.
+Moreover, $q$ can also be a new type of singularity, such as pseudo-singularity, or tangential singularity.
+See  Guardia et al [18].
 In this paper we only deal with crossing points.
@@
 We say that $q\in\Sigma$ is a crossing point of $\mathcal{X}$ if
-(2) $X^{+}(q),\nabla h(q)\right>\cdot\left<X^{-}(q),\nabla h(q)\right>>0,$
-[…]
-As far as we know the first approach on the characterization of global asymptotic stability for piecewise smooth vector fields is due Freire to et al [14], where they proved that the continuous matching of two Hurwitz vector fields is GAS.
+(2) $\left<X^{+}(q),\nabla h(q)\right>\cdot\left<X^{-}(q),\nabla h(q)\right>>0,$
+where $\left$$\cdot$$\cdot\right$> denotes the standard inner product of $\mathbb{R}^{2}$.
+Geometrically (2) means that $X^{+}(q)$ and $X^{-}(q)$ are transversal to $\Sigma$ at $q$ and both point towards either $\Sigma^{+}$ or $\Sigma^{-}$.
+In particular, the local trajectories of $X^{\pm}$ at $q$ agree on orientation and thus we can define the local trajectory of $\mathcal{X}$ at $q$ as the concatenation to the local trajectories of $X^{\pm}$.
+As far as we know the first approach on the characterization of global asymptotic stability for piecewise smooth vector fields is due to Freire et al [14], where they proved that the continuous matching of two Hurwitz vector fields is GAS.
 More precisely, if $\mathcal{X}=(X^{+},X^{-};\Sigma)$ is a planar piecewise linear vector field satisfying the following hypotheses:
@@
 imply the global asymptotically stability for $\mathcal{X}$.
+We remark that to know whether a piecewise linear system (and in particular piecewise Hurwitz system) is GAS has applicability in control theory.
+See Iwatani and Hara [21].
 In this paper we generalize these results in a different way.
@@
 Figure 1. Illustration of an orbit of the bouncing ball model. For simplicity, we interchanged the coordinate axes.
-At this instant, the ball bounces back upwards and its velocity undergoes an instantaneous change $v(t_{0})\mapsto-\rho v(t_{0})$, modeled by an inelastic collision with dissipate factor $10<\rho<$ 1.
+At this instant, the ball bounces back upwards and its velocity undergoes an instantaneous change $v(t_{0})\mapsto-\rho v(t_{0})$, modeled by an inelastic collision with dissipate factor $0<\rho<1$.
 That is, at $h=0$ our system undergoes a jump (or reset) $\varphi$ given by $\varphi(0,v)=(0,-\rho v)$.
@@
 Observe that any orbit converges to the origin, which represents a ball standing still in the ground.
-For more details on the bouncing ball model, we refer to [30, Section $2.2.3$] and [6, Section $1.2$].
-For a different example modeling a type of pinball machine with curved surface, we refer to [29, Example $2$].
+For more details on the bouncing ball model, we refer to [32, Section $2.2.3$] and [6, Section $1.2$].
+For a different example modeling a type of pinball machine with curved surface, we refer to [31, Example $2$].
 For a recent survey in the field, we refer to [5].
-The notion of a hybrid dynamical system is relatively new. Following Schaft and Schumacher [30, Section $1.1$], we quote:
-[$\dots$] the area of hybrid systems is still in its infancy and a general theory of hybrid systems seems premature.
+The notion of a hybrid dynamical system is relatively new. Following Schaft and Schumacher [32, Section $1.1$], we quote:
+$\dots$] the area of hybrid systems is still in its infancy and a general theory of hybrid systems seems premature.
 More inherently, hybrid systems is such a wide notion that sticking to a single definition shall be too restrictive.
-[$\dots$] Another difficulty in discussing hybrid systems is that various scientific communities with their own approaches have contributed (and are still contributing) to the area.
+$\dots$] Another difficulty in discussing hybrid systems is that various scientific communities with their own approaches have contributed (and are still contributing) to the area.
 At least the following three communities can be distinguished.
-[The] computer science community [$\dots$] [the] modeling and simulation community [$\dots$] [and the] systems and control community.
+[The] computer science community $\dots$] [the] modeling and simulation community $\dots$] [and the] systems and control community.
 Nevertheless, the field is expanding and in particular it has been receiving more attention in recent years from the community of Qualitative Theory of Differential Equations and Dynamical Systems.
@@
 However it follows from [22, 25] that if we embed such systems in a hybrid structure, then we have the existence of at least one limit cycle.
-As anticipated in the Introduction, in this paper we will embedded the the piecewise smooth vector field $\mathcal{X}=(X^{+},X^{-};\Sigma$ in a hybrid structure.
+As anticipated in the Introduction, in this paper we will embedded the the piecewise smooth vector field $\mathcal{X}=(X^{+},X^{-};\Sigma)$ in a hybrid structure.
 This embedding will allow us to draw a precise sufficient and necessary condition for global stability and moreover will point the way for the bifurcation of a limit cycle that would not be possible without the hybrid framework.
@@
 We shall also suppose that $\mathfrak{X}$ has the crossing property, i.e. we suppose that
-(5) $X^{+}(q),\nabla h_{\rho}(q)\right>\cdot\left<X^{-}(q),\nabla h_{\rho}(q)\right>>0,$
+(5) $\left<X^{+}(q),\nabla h_{\rho}(q)\right>\cdot\left<X^{-}(q),\nabla h_{\rho}(q)\right>>0,$
 for every $q\in\Sigma_{\rho}\setminus\{(0,0)\}$.
 In particular we recall that the geometric interpretation of (5) is that $X^{\pm}(q)$ are transversal to $\Sigma$ at $q$ and that both vectors $X^{+}(q)$ and $X^{-}(q)$ points to the same direction relatively to $\nabla h_{\rho}(q)$.
-For hybrid systems without the crossing property and thus allowing other dynamics such as sliding (or sticking), we refer to [30, Section $6.2.3$] and [6, p.
+For hybrid systems without the crossing property and thus allowing other dynamics such as sliding (or sticking), we refer to [32, Section $6.2.3$] and [6, p.
 $82$].
 The jump $\varphi=\varphi_{\rho}\colon\Sigma_{\rho}\to\Sigma_{\rho}$ is given by
-(6) $\varphi_{\rho}(x,y)=\left\{\begin{array}[]{ll}(-a|x|^{r},0),&\text{if }x\leqslant 0,\vspace{0.2cm}\\ (bx^{s},\rho bx^{s}),&\text{if }x\geqslant 0,\end{array}\right.$
+(6) $\varphi_{\rho}(x,y)=\left\{\begin{array}[]{ll}(-a|x|^{r},0),&\text{if }x\leqslant 0,\vskip 5.69046pt\\ (bx^{s},\rho bx^{s}),&\text{if }x\geqslant 0,\end{array}\right$.
 with $a$, $b$, $r$, $s\in\mathbb{R}_{>0}$. Observe that $\varphi_{\rho}$ has inverse given by
-(7) $\varphi_{\rho}^{-1}(x,y)=\left\{\begin{array}[]{ll}\bigl{(}-\bigl{|}a^{-1}x\bigr{|}^{\frac{1}{r}},0\bigr{)},&\text{if }x\leqslant 0,\vspace{0.2cm}\\ \bigl{(}b^{-\frac{1}{s}}x^{\frac{1}{s}},\rho b^{-\frac{1}{s}}x^{\frac{1}{s}}\bigr{)},&\text{if }x\geqslant 0,\end{array}\right.$
+(7) $\varphi_{\rho}^{-1}(x,y)=\left\{\begin{array}[]{ll}\bigl(-\bigl|a^{-1}x\bigr|^{\frac{1}{r}},0\bigr),&\text{if }x\leqslant 0,\vskip 5.69046pt\\ \bigl(b^{-\frac{1}{s}}x^{\frac{1}{s}},\rho b^{-\frac{1}{s}}x^{\frac{1}{s}}\bigr),&\text{if }x\geqslant 0,\end{array}\right$.
 for each $\rho\in\mathbb{R}_{\rho\geqslant 0}$.
@@
 Figure 2. Illustration of $\varphi_{\rho}$ and $\mathfrak{X}$. The origin $\mathcal{O}$ is interpreted as a singularity. By abuse of notation we drew two copies of $\Sigma_{\rho}$.
-The dynamics of $\mathfrak{X}=(X^{+},X^{-};\Sigma;\varphi)$ works as follows.
+The dynamics of $\mathfrak{X}=(X^{+},X^{-};\Sigma;\varphi)$ work as follows.
 Given $q_{0}\in\mathbb{R}^{2}\setminus\Sigma$, the local trajectory of $\mathfrak{X}$ at $q_{0}$ is given by the local trajectory of $X^{\pm}$.
@@
 If we let
-$B^{\pm}=\left(\begin{array}[]{cc}b_{11}^{\pm}&b_{12}^{\pm}\vspace{0.2cm}\\ b_{21}^{\pm}&b_{22}^{\pm}\end{array}\right),$
+$B^{\pm}=\left(\begin{array}[]{cc}b_{11}^{\pm}&b_{12}^{\pm}\vskip 5.69046pt\\ b_{21}^{\pm}&b_{22}^{\pm}\end{array}\right),$
 then the crossing property (5) restricted at $\Sigma^{2}_{\rho}$ writes
@@
 Consider the matrices
-$C^{\pm}=\left(\begin{array}[]{cc}c_{11}^{\pm}&c_{12}^{\pm}\vspace{0.2cm}\\ c_{21}^{\pm}&c_{22}^{\pm}\end{array}\right),$
+$C^{\pm}=\left(\begin{array}[]{cc}c_{11}^{\pm}&c_{12}^{\pm}\vskip 5.69046pt\\ c_{21}^{\pm}&c_{22}^{\pm}\end{array}\right),$
 where
-$\begin{array}[]{cc}\displaystyle c_{11}^{\pm}=1-\frac{(\delta^{\pm}-b_{12}^{\pm})\rho+b_{22}}{\eta^{\pm}}\rho,&\displaystyle c_{12}^{\pm}=\frac{(\delta^{\pm}-b_{12}^{\pm})\rho+b_{22}}{\eta^{\pm}},\vspace{0.2cm}\\ \displaystyle c_{21}^{\pm}=\rho+\frac{b_{12}^{\pm}\rho^{2}+b_{11}^{\pm}\rho-1}{\eta^{\pm}}\rho,&\displaystyle c_{22}^{\pm}=-\frac{b_{12}^{\pm}\rho^{2}+b_{11}^{\pm}\rho-1}{\eta^{\pm}}.\end{array}$
+$\begin{array}[]{cc}\displaystyle c_{11}^{\pm}=1-\frac{(\delta^{\pm}-b_{12}^{\pm})\rho+b_{22}}{\eta^{\pm}}\rho,&\displaystyle c_{12}^{\pm}=\frac{(\delta^{\pm}-b_{12}^{\pm})\rho+b_{22}}{\eta^{\pm}},\vskip 5.69046pt\\ \displaystyle c_{21}^{\pm}=\rho+\frac{b_{12}^{\pm}\rho^{2}+b_{11}^{\pm}\rho-1}{\eta^{\pm}}\rho,&\displaystyle c_{22}^{\pm}=-\frac{b_{12}^{\pm}\rho^{2}+b_{11}^{\pm}\rho-1}{\eta^{\pm}}.\end{array}$
 It follows from (8) that $C^{\pm}$ are well defined. Moreover we observe that
@@
 Therefore if we define the map $H\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$ by
-$H(x,y)=\left\{\begin{array}[]{ll}C^{+}\left(\begin{array}[]{c}x\\ y\end{array}\right),&\text{if }(x,y)\in\Sigma^{+},\vspace{0.2cm}\\ C^{-}\left(\begin{array}[]{c}x\\ y\end{array}\right),&\text{if }(x,y)\in\Sigma^{-},\end{array}\right.$
+$H(x,y)=\left\{\begin{array}[]{ll}C^{+}\left(\begin{array}[]{c}x\\ y\end{array}\right),&\text{if }(x,y)\in\Sigma^{+},\vskip 5.69046pt\\ C^{-}\left(\begin{array}[]{c}x\\ y\end{array}\right),&\text{if }(x,y)\in\Sigma^{-},\end{array}\right$.
 then it is a well defined topological equivalence between the piecewise systems $(X^{+},X^{-};\Sigma_{\rho})$ and $(Y^{+},Y^{-};\Sigma_{\rho})$. ∎
@@
 In particular, observe that $r_{2}^{+}<r_{1}^{+}<0$. Following [9] we define the sets
-$\begin{array}[]{l}R_{1}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x>r_{1}^{+}y\},\vspace{0.2cm}\\ E_{1}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x=r_{1}^{+}y\},\vspace{0.2cm}\\ R_{2}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon r_{2}^{+}y<x<r_{1}^{+}y\},\vspace{0.2cm}\\ E_{2}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x=r_{2}^{+}y\},\vspace{0.2cm}\\ R_{3}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x<r_{2}^{+}y\},\end{array}$
+$\begin{array}[]{l}R_{1}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x>r_{1}^{+}y\},\vskip 5.69046pt\\ E_{1}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x=r_{1}^{+}y\},\vskip 5.69046pt\\ R_{2}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon r_{2}^{+}y<x<r_{1}^{+}y\},\vskip 5.69046pt\\ E_{2}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x=r_{2}^{+}y\},\vskip 5.69046pt\\ R_{3}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x<r_{2}^{+}y\},\end{array}$
 see Figure 3.
@@
 Following [9] we define the sets
-$\begin{array}[]{l}S_{1}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x>r^{+}y\},\vspace{0.2cm}\\ E^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x=r^{+}y\},\vspace{0.2cm}\\ S_{2}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x<r^{+}y\},\end{array}$
+$\begin{array}[]{l}S_{1}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x>r^{+}y\},\vskip 5.69046pt\\ E^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x=r^{+}y\},\vskip 5.69046pt\\ S_{2}^{+}=\{(x,y)\in\Sigma_{\rho}^{+}\setminus\Sigma_{\rho}\colon x<r^{+}y\},\end{array}$
 see Figure 4.
@@
     If $q\in S_{2}^{+}$, then there is $t^{2}_{q}>0$ such that $X^{+}(t^{2}_{q},q)\in\Sigma_{\rho}$.
-4.4. Hurwitz vector fields of type $F$
+[…]
 Suppose that the origin is an attracting focus of $X^{+}$ with eigenvalues given by
@@
 For interpretation of the references to color in this figure legend, the reader is referred to the web version of this article.
-To end this section, we observe that since $X^{+}$ is linear we can compute explicitly its solution, see [28, Chapter $1$].
+To end this section, we observe that since $X^{+}$ is linear we can compute explicitly its solution, see [30, Chapter 11].
 Therefore it is not hard to see that if $X^{+}$ is a Hurwitz vector field of type $F$ with the normal form given by Lemma 1, then its solution
@@
 is given by
-(9) $\begin{array}[]{l}\displaystyle x^{+}(t;x,y)=e^{\lambda^{+}t}\left(x\cos(\mu^{+}t)+\frac{1}{\mu^{+}}(\lambda^{+}x-r_{1}^{+}r_{2}^{+}y)\sin(\mu^{+}t)\right),\vspace{0.2cm}\\ \displaystyle y^{+}(t;x,y)=e^{\lambda^{+}t}\left(y\cos(\mu^{+}t)+\frac{1}{\mu^{+}}(x-\lambda^{+}y)\sin(\mu^{+}t)\right).\end{array}$
+(9) $\begin{array}[]{l}\displaystyle x^{+}(t;x,y)=e^{\lambda^{+}t}\left(x\cos(\mu^{+}t)+\frac{1}{\mu^{+}}(\lambda^{+}x-r_{1}^{+}r_{2}^{+}y)\sin(\mu^{+}t)\right),\vskip 5.69046pt\\ \displaystyle y^{+}(t;x,y)=e^{\lambda^{+}t}\left(y\cos(\mu^{+}t)+\frac{1}{\mu^{+}}(x-\lambda^{+}y)\sin(\mu^{+}t)\right).\end{array}$
 5. Proof of Theorem A
@@
 From (14) we can deduce statements $(i)$, $(ii)$, $(iii)$ and $(iv)$.
-More precisely, if $r=1/s$ then we can factor $x$ out of (14) and thus we obtain that the origin is globally asymptotically stable (resp.
+More precisely, if $r=1/s$ then we can factor $x$ out of (14) and thus obtain that the origin is globally asymptotically stable (resp.
 globally asymptotically unstable) if (14) is satisfied (resp.
@@
 regardless of $X^{\pm}$.
-This in addition with Theorem A$(a)$ ensures that if the jump $\varphi_{\rho}$ is reduced to the identity map (i.e.
+This in addition with Theorem A(a)(a) ensures that if the jump $\varphi_{\rho}$ is reduced to the identity map (i.e.
 if the hybrid system $\mathfrak{X}$ is reduced to a piecewise linear vector field) and $\Sigma_{\rho}$ is reduced to a straight line, then $\mathfrak{X}$ is GAS.
@@
 Remark 2 (The limit case $\rho\to+\infty$).
-From (15) we have that,
+From (15) we have,
 $\lim\limits_{\rho\to+\infty}t^{\pm}(\rho)=\frac{1}{\mu^{\pm}}\left(\pm\arctan\left(\frac{\mu^{\pm}}{\lambda^{\pm}}\right)+\pi\right).$
 This in addition with the following well-known properties of the trigonometric functions
-$\begin{array}[]{ll}\displaystyle\cos(x+\pi)=-\cos x,&\displaystyle\sin(x+\pi)=-\sin(x),\vspace{0.2cm}\\ \displaystyle\cos(\arctan x)=\frac{1}{\sqrt{1+x^{2}}},&\displaystyle\sin(\arctan x)=\frac{x}{\sqrt{1+x^{2}}},\end{array}$
+$\begin{array}[]{ll}\displaystyle\cos(x+\pi)=-\cos x,&\displaystyle\sin(x+\pi)=-\sin(x),\vskip 5.69046pt\\ \displaystyle\cos(\arctan x)=\frac{1}{\sqrt{1+x^{2}}},&\displaystyle\sin(\arctan x)=\frac{x}{\sqrt{1+x^{2}}},\end{array}$
 and some straightforward calculations, allow us to conclude that
-(16) $\lim\limits_{\rho\to+\infty}\frac{|\cos(\mu^{-}t^{-})-\Phi^{-}\sin(\mu^{-}t^{-})|}{|\cos(\mu^{+}t^{+})+\Phi^{+}\sin(\mu^{+}t^{+})|^{r}}=\left\{\begin{array}[]{ll}0,&\text{if }r>1,\vspace{0.2cm}\\ +\infty,&\text{if }r<1,\vspace{0.2cm}\\ \displaystyle\sqrt{\frac{(\lambda^{-})^{2}+(\mu^{-})^{2}}{(\lambda^{+})^{2}+(\mu^{+})^{2}}},&\text{if }r=1.\end{array}\right.$
+(16) $\lim\limits_{\rho\to+\infty}\frac{|\cos(\mu^{-}t^{-})-\Phi^{-}\sin(\mu^{-}t^{-})|}{|\cos(\mu^{+}t^{+})+\Phi^{+}\sin(\mu^{+}t^{+})|^{r}}=\left\{\begin{array}[]{ll}0,&\text{if }r>1,\vskip 5.69046pt\\ +\infty,&\text{if }r<1,\vskip 5.69046pt\\ \displaystyle\sqrt{\frac{(\lambda^{-})^{2}+(\mu^{-})^{2}}{(\lambda^{+})^{2}+(\mu^{+})^{2}}},&\text{if }r=1.\end{array}\right$.
 Since the above limit is the only factor of (14) that depends on $\rho$, we conclude that if $r\neq 1$ then the right-hand side of (14) can either explode or collapse to zero as $\rho\to+\infty$.
@@
 in the proof of Theorem A. Using the global normal form provided by Lemma 1, we have that $\mathfrak{X}$ is given by
-$\begin{array}[]{l}X^{\pm}(x,y)=(-2x-2y,x),\vspace{0.2cm}\\ \Sigma=\{(x,y)\in\mathbb{R}^{2}\colon y=0\},\vspace{0.2cm}\\ \varphi(x,0)=(x^{3},0).\end{array}$
+$\begin{array}[]{l}X^{\pm}(x,y)=(-2x-2y,x),\vskip 5.69046pt\\ \Sigma=\{(x,y)\in\mathbb{R}^{2}\colon y=0\},\vskip 5.69046pt\\ \varphi(x,0)=(x^{3},0).\end{array}$
 Replacing (17) into (12), (13) and (15) we obtain
@@
 Acknowledgments
-The first author is partially supported by Fundação de Amparo à Pesquisa do Estado de Minas Gerais, Brazil, grant APQ-02153-23; by São Paulo Research Foundation (FAPESP), Brazil, grant 2019/07316–0; and by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Brazil, grant 311921/2020-5.
+The first author is partially supported by Fundação de Amparo à Pesquisa do Estado de Minas Gerais, Brazil, grant numbers APQ-02153-23, APQ-05207-23 and RED-00133-21.
 The second author is partially supported by São Paulo Research Foundation (FAPESP), Brazil, grant 2021/01799-9.